OpenStudy (babynini):

Find two polar coordinate representations for the rectangular coordinate point (-6, 2 (sqroot3)), one with r>0, one with r<0 and both with

OpenStudy (babynini):

For r I already got 4 (sqroot3) and -4 (Sqroot3)

OpenStudy (babynini):

and then tan(theta)= -(2(sqroot3))/6

OpenStudy (babynini):

@rvc :) I know you just got on. But if you're free!

rvc (rvc):

ah bit busy dear :( @JoannaBlackwelder

OpenStudy (babynini):

\[\tan(\theta) = -\frac{ 2\sqrt{3} }{ 6 }\]

OpenStudy (joannablackwelder):

The tan equation looks good to me :-) Can you solve for theta?

OpenStudy (babynini):

em theta = arctan (x) but i'm not sure how to keep it not in decimal form

OpenStudy (babynini):

....?

OpenStudy (joannablackwelder):

We will need to use the unit circle and the idea that tan is sin/cos

OpenStudy (joannablackwelder):

http://www.mathsisfun.com/geometry/unit-circle.html Look down at the very bottom chart.

OpenStudy (babynini):

I'm not finding anywhere that tangent equals that on the unit circle. It is sq3/3 at pi/6

OpenStudy (babynini):

and -sq3/3 at 11pi/6

OpenStudy (joannablackwelder):

Right, and -2sqrt3/6 is -sqrt3/3

OpenStudy (babynini):

oh.. We can take the 6 and divide it by the 2 in the numerator?

OpenStudy (joannablackwelder):

Yep, those can simplify, since they have a common factor of 2

OpenStudy (joannablackwelder):

Do you see another one other than 11pi/6?

OpenStudy (babynini):

5pi/6

OpenStudy (babynini):

so how do I know which to choose? 11pi/6 or 5pi/6 ?

OpenStudy (joannablackwelder):

Awesome! We are looking for 2 options, so both.

OpenStudy (joannablackwelder):

And I don't get what you got for r. How did you get that?

OpenStudy (babynini):

Give me a moment to type that up :)

OpenStudy (joannablackwelder):

Oh, sorry, my mistake. You're rs are correct! :-)

OpenStudy (joannablackwelder):

As long as you can combine them correctly.

OpenStudy (joannablackwelder):

Into two pairs.

OpenStudy (babynini):

\[r^2=x^2+y^2\] \[r^2=-6^2+(2\sqrt{3})^2\] \[r^2=36+12 = \sqrt{48}\] Simplify and get \[r=4\sqrt{3}\]

OpenStudy (babynini):

Right, how do I know how to pair them?

OpenStudy (babynini):

(4 sq3, 5pi/6) (-4 sq3, 11pi/6) right? but i'm not sure how I got there xD or why

OpenStudy (joannablackwelder):

Let's look at each of those graphically to see if they look like they are in the right spots.

OpenStudy (joannablackwelder):